prove that the path followed by a projectile is a straigt line as seen from another projectile.

Grade:11

3 Answers

vikas askiitian expert
13 years ago

eq of path of projectile is s=ucos@ -gt2/2

if two projectiles are projected from horizontal ground with different initial velocity and angle....both these quantities are known then at any time t

coordinates of A,  S1=(UacosA)t i - gt2/2 j

coordinates of B, S2=(UbcosB)t i - gt2/2j

 coordinates of A wrt to B is S2-S1=t(UacosA-UbcosB)i + 0 j

            Xi + Yj = kt i + 0j                                     (UacosA-UbcosB = constant=k)

  ie X=kt  is a straight line              

Anuraj
6 years ago
if two projectiles are projected from horizontal ground with different initial velocity and angle....both these quantities are known then at any time t coordinates of A, S1=(UacosA)t i - gt2/2 j coordinates of B, S2=(UbcosB)t i - gt2/2j coordinates of A wrt to B is S2-S1=t(UacosA-UbcosB)i + 0 j Xi + Yj = kt i + 0j (UacosA-UbcosB = constant=k) ie X=kt is a straight line
As
5 years ago
X1 = u1cosa * t 
X2 = u2cosb*t
Y2= u2sinb *t - 1/2 gt*t
Y1= u1sina*t -1/2 gt*t
Let Y2 -Y1 = Y
Let X2 -X1 = X
Then X/Y = u2cosb - u1cosa / u2sinb - u1sina
X/Y = constant
X/Y = k
X = kY
Hence it is a equation of straight line